### 20 Mar 2024 v1

# Enhanced Cauchy Schwarz inequality and some of its statistical applications

Abstract
We present a general refinement of the Cauchy-Schwarz inequality over complete inner product spaces and show that it can be of interest for some statistical applications. This generalizes and simplifies previous results on the same subject.
keywords:
Cauchy-Schwarz inequality, Orthogonal projections, Correlation indexes, Cauchy-Schwarz divergence, Nonparametric estimators
1 Introduction
It is widely recognized that Cauchy-Schwarz (CS) inequality is one of the most important inequality in mathematical analysis, if not the most important one, see Steele (2004) and Dragomir (2003). This reputation comes from its large spectrum of applications in different fields of science ranging from physics to probability and statistics, but including also signal analysis, coding theory and computer science. It is ascertained that the Russian mathematician V. Bunyakovsky has contributed to its discovery and diffusion. It is also well known that the basic CS inequality admits several generalizations to function, matrix and operator spaces, perhaps what is less known is that the basic inequality itself admits improvements. The aim of the present paper is to show that there are cases where the usage of a refined form of the CS inequality turns out to be easy to apply and useful to consider. We shall show that fact by a set of examples arising in the fields of statistics and data analysis.
2 Enhanced Cauchy-Schwarz inequality
Let $(H,<,>)$ a real or complex Hilbert space and $||x||=\sqrt{<x,x>}$ the vector norm, the celebrated Cauchy-Schwarz inequality states that
$|<x,y>|\leq||x||\cdot||y||\;\;\;\forall x,y\in H.$ (1)
In what follows only real Hilbert spaces will be contemplated. In a finite dimensional setting let us take $x,y\in H=\mathbb{R}^{n}$, endowed with the Euclidean norm, and consider the anti-symmetric square matrix $C=C(x,y)$ having elements defined as $(c_{ij})\equiv\frac{1}{\sqrt{2}}(x_{i}y_{j}-x_{j}y_{i})$, $i,j=1,\ldots,n$, with $||C||_{2}\equiv\sqrt{\sum_{i,j}c^{2}_{ij}}$ its $L^{2}$-norm. Then CS inequality stems from the evaluation of the matrix norm, that is:
$||C||_{2}^{2}=||x||^{2}||y||^{2}-<x,y>^{2},$ (2)
an identity which can be traced back to Lagrange, see Steele (2004), and Magnus, Neudecker (2019). The following proposition, which sharpens CS inequality, is a consequence of the existence of orthogonal projections on closed subspaces of a Hilbert space, a classical theorem, see Berberian (1961). We state it in the following useful form which we could not find in the literature:
Proposition 2.1: Let $(H,<\cdot,\cdot>)$ a real Hilbert space, $V$ any closed subspace of $H$. Let $P\equiv P_{V}$ denote the orthogonal projection on $V$, that is $PH=V$, and define the following symmetric function:
$D(x,y|P)\equiv||Px||\cdot||Py||+||x-Px||\cdot||y-Py||\;\;\forall x,y\in H.$ (3)
Then it holds
$|<x,y>|\leq D(x,y|P)\leq||x||\cdot||y||\;\;\forall x,y\in H.$ (4)
Furthermore, let $\mathcal{P}$ denote the family of all orthogonal projections relative to all possible closed subspaces of $H$, then for each pair $(x,y)$ we have
$D^{+}(x,y)\equiv\sup_{P\in\mathcal{P}}D(x,y|P)=||x||\cdot||y||$ (5)
and
$D^{-}(x,y)\equiv\inf_{P\in\mathcal{P}}D(x,y|P)=|<x,y>|.$ (6)
The enhanced CS inequality is therefore the statement that for any orthogonal projection $P$ it holds
$D^{-}(x,y)\leq D(x,y|P)\leq D^{+}(x,y),\;\;\forall x,y\in H.$ (7)
Remark2.1: Notice the following three properties of the symmetric function $D(\cdot,\cdot|P)$ : (i) $D(x,x|P)=||x||^{2}$; (ii) $D(Px,y|P)=||Px||\;||Py||$; (iii) $D(x,y|P)=0\;\text{for all}\;(x,y)\in V\times V^{\perp}$, with $V=PH$.
proof: Writing $x=Px+(I-P)x$ by orthogonality we have $||x||^{2}=p(x)^{2}+q(x)^{2}$, where $p(x)\equiv||Px||$ and $q(x)\equiv\sqrt{||x||^{2}-p(x)^{2}}=||(I-P)x||$; therefore
$|<x,y>|=|<Px,Py>+<(I-P)x,(I-P)y>|\leq|<Px,Py>|+|<(I-P)x,(I-P)y>|$
$\leq p(x)p(y)+q(x)q(y)\leq\sqrt{p(x)^{2}+q(x)^{2}}\sqrt{p(y)^{2}+q(y)^{2}}=||x||\cdot||y||$
by applying orthogonality, triangle inequality, the Cauchy-Schwarz inequality and noticing
$(p(x)p(y)+q(x)q(y))^{2}+(p(x)q(y)-p(y)q(x))^{2}=(p(x)^{2}+q(x)^{2})(p(y)^{2}+q(y)^{2}).$ (8)
The second part of the statement follows by choosing $V=\{0\}$ or $V=H$, that is $PH=\{0\}$ or $PH=H$ in order to get $D^{+}(x,y)$. On the other hand the choice of $V=\{\lambda x\}$, that is of $P=P_{x}$, with $P_{x}z\equiv<{\frac{x}{||x||}},z>\frac{x}{||x||}$ leads to $D^{-}(x,y)$; indeed
$D(x,y|P_{x})=||P_{x}x||\cdot||P_{x}y||+||x-P_{x}x||\cdot||y-P_{x}y||$
$=||x||\cdot||P_{x}y||=||x||\cdot|<{\frac{x}{||x||}},y>|=|<x,y>|.$
Corollary 2.1 (triangle inequality): Let $(X,<\cdot,\cdot>)$ a real Hilbert. Then for any orthogonal projection $P$ on a closed subspace we have
$||x+y||\leq\sqrt{(||Px||+||Py||)^{2}+(||(I-P)x||+||(I-P)y||)^{2}}\leq||x||+||y||\;\;\forall x,y\in X.$ (9)
To grasp the content of Proposition 2.1 let us consider the following simple but instructive example:
Example2.1 Let $H=\mathbb{R}^{n}$, endowed with the Euclidean distance $d_{n}$; for $k\geq 1$ consider the subspace $V_{1:k}\equiv\{z|z=(z_{1},z_{2},\ldots,z_{k},0,\ldots,0)\}$ and its orthogonal complement $V^{\perp}_{1:k}=V_{k+1:n}\equiv\{z|z=(0,\ldots,0,z_{k+1},z_{k+2},\ldots,z_{n})\}$ Define the orthogonal projection $P_{k}H\equiv V_{1:k}$, then we have
$D(x,y|P_{k})=\sqrt{x_{1}^{2}+\ldots+x^{2}_{k}}\sqrt{y_{1}^{2}+\ldots+y_{k}^{2}}+\sqrt{x_{k+1}^{2}+\ldots+x^{2}_{n}}\sqrt{y_{k+1}^{2}+\ldots+y_{n}^{2}}$ (10)
and by Proposition 2.1 the following inequalities
$|\sum_{i=1}^{n}x_{i}y_{i}|\leq D(x,y|P_{k})\leq\sqrt{x_{1}^{2}+\ldots+x_{n}^{2}}\cdot\sqrt{y_{1}^{2}+\ldots y^{2}_{n}}\;\;\;\forall x,y\in\mathbb{R}^{n}$ (11)
hold, with matrix form of $P_{k}$ w.r.t the canonical basis of $\mathbb{R}^{n}$ given by $diag(I_{k\times k},{\bm{0}}_{(n-k)\times(n-k)})$.
The interpretation of (11) is straightforward: pick two arbitrary points A and B in the $n$-dimensional euclidean space and orthogonally project them on $V_{1:k}$ getting points $\tilde{A}$ and $\tilde{B}$ and on $V^{\perp}_{1:k}$ getting points $\hat{A}$ and $\hat{B}$. Then the number $d_{n}(0,\tilde{A})d_{n}(0,\tilde{B})+d_{n}(0,\hat{A})d_{n}(0,\hat{B})$ is not greater than the number $d_{n}(0,A)d_{n}(0,B)$ and not smaller than the value of the scalar product of the two vectors pointing from the origin to $A$ and $B$.
3 Statistical applications of the enhanced CS inequality
In this section, by means of a series of different examples, we aim to show that sharpened CS inequalities may find applications in applied statistics. The underlying idea being that, in relation to the problem under investigation, a suitable choice of the relevant subspace and of the associated orthogonal projection $P$ can produce a better estimation of the quantities of interest. In the first example the chosen subspace will have dimension one.
Example3.1 (Sample covariance) Let $H=\mathbb{R}^{n}$, $e=(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}})$ and $u=\sqrt{n}e$. Define $Px\equiv<x,e>e$ for all $x\in H$, by Proposition 2.1 we have
$|\sum_{i=1}^{n}x_{i}y_{i}|\leq|<x,e>|\cdot|<y,e>|+\sqrt{||x||^{2}-|<x,e>|^{2}}\cdot\sqrt{||y||^{2}-|<y,e>|^{2}}$
$=\frac{|<x,u>|}{\sqrt{n}}\cdot\frac{|<y,u>|}{\sqrt{n}}+\sqrt{||x||^{2}-\biggl{(}\frac{|<x,u>|}{\sqrt{n}}\biggr{)}^{2}}\cdot\sqrt{||y||^{2}-\biggl{(}\frac{|<y,u>|}{\sqrt{n}}\biggr{)}^{2}}.$
therefore by setting $\bar{x}=\frac{x_{1}+\ldots+x_{n}}{n}$ we obtain the estimate:
$|\sum_{i=1}^{n}x_{i}y_{i}|\leq n|\bar{x}\bar{y}|+\sqrt{||x||^{2}-n|\bar{x}|^{2}}\cdot\sqrt{||y||^{2}-n|\bar{y}|^{2}}.$ (12)
Notice that by applying (12) to $x^{\prime}=x-\bar{x}u$ and $y^{\prime}=y-\bar{y}u$ we get the sample covariance estimate
$\big{|}<x,y>-n\;\bar{x}\bar{y}\big{|}\leq||x^{\prime}||\;||y^{\prime}||=n\cdot{\overline{s}(x)}{\overline{s}(y)},$ (13)
where $\overline{s}(x)^{2}\equiv\frac{1}{n}||x||^{2}-\overline{x}^{2}$. Furthermore, squaring (12) gives
$(\sum_{i=1}^{n}x_{i}y_{i})^{2}\leq\biggl{(}\;n|\bar{x}\bar{y}|+\sqrt{||x||^{2}-n|\bar{x}|^{2}}\cdot\sqrt{||y||^{2}-n|\bar{y}|^{2}}\;\biggr{)}^{2}=$ (14)
$=||x||^{2}||y||^{2}-n\biggr{(}|\bar{x}|\sqrt{||y||^{2}-n|\bar{y}|^{2}}-|\bar{y}|\sqrt{||x||^{2}-n|\bar{x}|^{2}}\biggl{)}^{2}.$
Finally we notice that, in the same vein, one may choose $H=L^{2}(\mathcal{F},Q)$, $\mathcal{F}$ a sigma-algebra, $Q$ a probability measure, and $<X,Y>\equiv\mathbb{E}(XY)$. For any $X\in H$ define the one-dimensional projection $PX\equiv<X,1>1=(\mathbb{E}X)1$. In this case $D(X,Y|P)=|\mathbb{E}X||\mathbb{E}Y|+\sigma_{X}\sigma_{Y}$ and (8) implies
$(\mathbb{E}(XY))^{2}\leq||X||^{2}||Y||^{2}-\biggr{(}|\mathbb{E}X|\sigma_{Y}-|\mathbb{E}Y|\sigma_{X}\biggl{)}^{2}\leq||X||^{2}||Y||^{2}\;\;\forall X,Y\in H,$ (15)
which is the main result presented in Walker (2017), jointly with a pair of nice applications. It must be remarked that in case both $X$ and $Y$ have zero mean the estimate (15) offers no improvement over CS, and the same it holds for centered data in the previously displayed sample estimates. This is due to the fact that the chosen $P$ is projecting onto a subspace orthogonal to $W=\{X\in L^{2}:\mathbb{E}(X)=0\}$.
Example3.2(Sample cross-covariance) Let us consider the sample cross-covariance function $\overline{R}_{(x,y)}(h)$ among two temporal sequences of observations $(x_{1},x_{2},\ldots,x_{n})$ and $(y_{1},y_{2},\ldots,y_{n})$:
$\overline{R}_{(x,y)}(h)\equiv\frac{1}{n}\sum_{t=1}^{n-h}(x_{t}-\bar{x})(y_{t+h}-\bar{y}),$
with $h=1,\ldots,n-1$. In the following we shall suppose the observed data being generated by two mean-zero ergodic processes so, by assuming $n$ sufficiently large, we may take $\bar{x}=\bar{y}=0$. For a fixed $h$ we take $H=\mathbb{R}^{n-h}$, and choose the projection $P_{k}$, $k\leq n-h$, as in Example 2.1; by using the first inequality in (11), we have
$|\overline{R}_{(x,y)}(h)|=\frac{1}{n}|x_{1}\underbrace{y_{1+h}}_{z_{1}}+x_{2}\underbrace{y_{2+h}}_{z_{2}}+\ldots+x_{n-h}\underbrace{y_{n}}_{z_{n-h}}|\leq\frac{1}{n}D(x,z|P_{k}),$ (16)
with
$D(x,z|P_{k})=\sqrt{x_{1}^{2}+\ldots+x^{2}_{k}}\sqrt{y_{1+h}^{2}+\ldots+y_{k+h}^{2}}+\sqrt{x_{k+1}^{2}+\ldots+x^{2}_{n-h}}\sqrt{y_{k+h+1}^{2}+\ldots+y_{n}^{2}}.$
For instance, setting $z_{r:s}\equiv(z_{r},\ldots,z_{s})\in\mathbb{R}^{s-r+1}$, $1\leq r\leq s\leq n$ and choosing in (16) $k=h$ (for $h=1,\ldots,[n/2]$),the dimension of the subspace equalizing the number of the forward time-shift steps, leads to
$|\overline{R}_{(x,y)}(h)|\leq(\frac{h}{n})\sqrt{\overline{x^{2}}_{1:h}}\sqrt{\overline{y^{2}}_{h+1:2h}}+(1-\frac{2h}{n})\sqrt{\overline{x^{2}}_{h+1:n-h}}\sqrt{\overline{y^{2}}_{2h+1:n}}\;,$
which is a linear combination of products of sample standard deviations on different time intervals improving over the classical CS inequality estimate.
Example3.3 (Correlation and conditioning) We take $H=L^{2}(\mathcal{F},Q)$ as in Example 3.1. By setting $X^{\prime}\equiv X-\mathbb{E}X$ and $Y^{\prime}\equiv Y-\mathbb{E}Y$, CS inequality implies
$|cov(X,Y)|=|\mathbb{E}(X^{\prime}Y^{\prime})|\leq||X^{\prime}||\;||Y^{\prime}||=\sigma_{X}\sigma_{Y},$ (17)
from which the classical bound $|\rho_{X,Y}|\leq 1$. Consider now any sub-sigma algebra $\mathcal{G}\subseteq\mathcal{F}$ and the closed subspace of $H$ given by $V_{\mathcal{G}}\equiv L^{2}(\mathcal{G},Q)$. Let $P_{\mathcal{G}}H\equiv V_{\mathcal{G}}$ the associated orthogonal projection, henceforth for $X\in H$, we have $P_{\mathcal{G}}X=\mathbb{E}(X|\mathcal{G})$. Hence $P_{\mathcal{G}}X^{\prime}=\mathbb{E}(X|\mathcal{G})-\mathbb{E}X$ and
$D(X^{\prime},Y^{\prime}|P_{\mathcal{G}})=||\mathbb{E}(X^{\prime}|\mathcal{G})||\cdot||\mathbb{E}(Y^{\prime}|\mathcal{G})||+||X^{\prime}-\mathbb{E}(X^{\prime}|\mathcal{G})||\cdot||Y^{\prime}-\mathbb{E}(Y^{\prime}|\mathcal{G})||$
$=||\mathbb{E}(X|\mathcal{G})-\mathbb{E}X||\cdot||\mathbb{E}(Y|\mathcal{G})-\mathbb{E}Y||+||X-\mathbb{E}(X|\mathcal{G})||\cdot||Y-\mathbb{E}(Y|\mathcal{G})||$
$=\sigma_{\mathbb{E}(X|\mathcal{G})}\;\sigma_{\mathbb{E}(Y|\mathcal{G})}+\sqrt{\sigma^{2}_{X}-\sigma^{2}_{\mathbb{E}(X|\mathcal{G})}}\;\sqrt{\sigma^{2}_{Y}-\sigma^{2}_{\mathbb{E}(Y|\mathcal{G})}}$
By Proposition 2.1 we have $|\mathbb{E}(X^{\prime}Y^{\prime})|\leq D(X^{\prime},Y^{\prime}|P_{\mathcal{G}})\leq||X^{\prime}||\cdot||Y^{\prime}||$ that is
$|cov(X,Y)|\leq\sigma_{\mathbb{E}(X|\mathcal{G})}\;\sigma_{\mathbb{E}(Y|\mathcal{G})}+\sqrt{\sigma^{2}_{X}-\sigma^{2}_{\mathbb{E}(X|\mathcal{G})}}\;\sqrt{\sigma^{2}_{Y}-\sigma^{2}_{\mathbb{E}(Y|\mathcal{G})}}\leq\sigma_{X}\sigma_{Y},$ (18)
For instance, in case $X\in\mathcal{G}$ and $Y\in\mathcal{G}$ then (18) reduces to (17), while in case $X\notin\mathcal{G}$ and $Y\in\mathcal{G}$ then (18) takes the form
$|cov(X,Y)|\leq\sigma_{\mathbb{E}(X|\mathcal{G})}\cdot\sigma_{Y}\leq\sigma_{X}\sigma_{Y}.$ (19)
Notice that when $\mathcal{G}\equiv\sigma(Y)$ and the variables $(X,Y)$ are jointly normal we have $\mathbb{E}(X|\mathcal{G})=\mu_{X}+\rho_{X,Y}\frac{\sigma_{X}}{\sigma_{Y}}(Y-\mu_{Y})$, so the lower inequality in (19) becomes an equality.
Finally, the above estimates suggest the consideration of correlation coefficients defined as
$\rho_{X,Y}^{\mathcal{G}}\equiv\frac{E(X^{\prime}Y^{\prime})}{D(X^{\prime},Y^{\prime}|P_{\mathcal{G}})},$ (20)
(in case $\rho_{X,Y}^{\mathcal{G}}=\frac{0}{0}$ its value is set to zero) which values depend on the sigma-algebra $\mathcal{G}$ inherent to the problem under investigation. Indeed it holds $|\rho_{X,Y}|\leq|\rho_{X,Y}^{\mathcal{G}}|\leq 1$, therefore encoding information might significantly enlarge the range of the estimated correlation. When $\mathcal{G}$ is the trivial sigma-algebra the two numbers $\rho_{X,Y}$ and $\rho_{X,Y}^{\mathcal{G}}$ coincide, and the random variable $\rho_{X,Y|\mathcal{G}}$, defined in the standard way, also equalizes this value.
Remark3.1:Let $X,Y\in L^{2}(\mathcal{F},Q)$, generalizing the idea of the previous example, we can use the lower inequality in (4) to define a P-correlation coefficient $\rho_{X,Y}(P)$ as
$\rho_{X,Y}(P)\equiv\frac{cov(X,Y)}{D(X,Y|P)}\in[-1,1],$ (21)
Notice that $\rho_{X,Y}(I)=\rho_{X,Y}$ but in general $\rho_{X,Y}(P)\neq\rho_{PX,PY}$; obviously $\rho_{X,Y}(P_{\mathcal{G}})=\rho_{X,Y}^{\mathcal{G}}$.
Example3.4 (CS divergence) Consider the Hilbert space $L^{2}\equiv L^{2}(\mathbb{R}^{d},dx)$ and $\text{Dens}^{2}\equiv L^{2}\cap\text{Dens}(\mathbb{R}^{d})$, where $\text{Dens}(\mathbb{R}^{d})$ is set of all probability density functions over $\mathbb{R}^{d}$. It is then possible to introduce the functional $\mathcal{D}iv(\cdot,\cdot|P):\text{Dens}^{2}\times\text{Dens}^{2}\to[0,\infty]$ defined as follows
[…]
Recalling that $D(f,f|P)=||f||^{2}$ we have $\mathcal{D}iv(f,g|P)=0$ for $f=g$. For $P=I$ the functional is known as the Cauchy-Schwarz divergence and has found application in density based clustering and machine learning, Jenssen et al.(2006), Kampffmeyer et al.(2019). CS divergence does not verify the triangle inequality and therefore does not define a true distance but only a pseudo-distance, however it is symmetric, positive and null on the diagonal. The CS P-divergence introduced by (22) shares exactly the same properties, moreover the following holds
$\mathcal{D}iv(f,g|I)-\mathcal{D}iv(f,g|P)=-log\biggl{(}\frac{D(f,g|P)}{||f||\;||g||}\biggr{)}\geq 0.$
Therefore, using (4), we have $\mathcal{D}iv(f,g|I)\geq\mathcal{D}iv(f,g|P)\geq 0$. Furthermore, if $(P_{N})_{N\geq 1}$ is such that $||P_{N}u||\to||u||$ $\forall u\in L^{2}$, then $D(f,g|P_{N})\to||f||\cdot||g||$ and the CS $P_{N}$-divergences converge to the CS divergence. This last property allows for a nonparametric estimation of the CS divergence among two densities based on projection estimators, see Tsybakov (2009), which we are going to outline fixing $d=1$. Let $(e_{k}(\cdot))_{k\geq 1}$ be an orthonormal basis, for any $u\in L^{2}$ set $u_{k}\equiv<u,e_{k}>$, so that $||u||^{2}=\sum_{k=1}^{\infty}u_{k}^{2}$. For $N\geq 1$ define orthogonal projections $P_{N}u\equiv\sum_{k=1}^{N}u_{k}e_{k}$, clearly $||P_{N}u||\to||u||$ $\forall u\in L^{2}$. Suppose $(X_{1},\ldots,X_{n})$ be i.i.d. random variables having density $f$ and $(Y_{1},\ldots,Y_{n})$ be i.i.d. random variables having density $g$ and define unbiased estimators for $f_{k}$ and $g_{k}$ respectively as $\hat{f}_{k,n}\equiv n^{-1}\sum_{i=1}^{n}e_{k}(X_{i})$ and $\hat{g}_{k,n}\equiv n^{-1}\sum_{j=1}^{n}e_{k}(Y_{j})$. It follows that $t_{N}(f)\equiv||P_{N}f||^{2}$ can be estimated by $\hat{t}_{N,n}(f)\equiv\sum_{k=1}^{N}\hat{f}_{k,n}^{2}$ and similarly $t_{N}(g)$ can be estimated by $\hat{t}_{N,n}(g)$. Henceforth, reconsideration of the basic Example 2.1 suggests that the CS $P_{N}$-divergence $T_{N}\equiv Div(f,g|P_{N})$ may be estimated by using the statistics
$\hat{T}_{N,n}({\mathbb{X}},\!{\mathbb{Y}})\equiv log\biggl{(}\frac{\sqrt{\hat{t}_{N,n}(f)}\sqrt{\hat{t}_{N,n}(g)}+\sqrt{\hat{r}_{N,n}(f)}\sqrt{\hat{r}_{N,n}(g)}}{\sum_{k=1}^{2N}\hat{f}_{k,n}\hat{g}_{k,n}}\biggr{)},$
where $\hat{r}_{N,n}(f)\equiv\sum_{k=N+1}^{2N}\hat{f}_{k,n}^{2}$, so that for $n$ and $N$ sufficiently large this leads to an estimation of the CS divergence of the two densities.
Declaration:This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.